Lisa will be half as old as her brother in five years. How old will she be in seven years if her brother is now nine?
Answer by lwsshak3(11628)

Lisa will be half as old as her brother in five years. How old will she be in seven years if her brother is now nine?
***
At present:
Brother's age=9
..
In 5 years:
Brother's age=9+5=14
Lisa's age=14/2=7
..
In 7 years:
Lisa's age=7+2=9